OpenStudy (anonymous):

PLEASE HELP!!!! At the concession standy, a variety of combos can be purchased. A combo includes a meat, a snack, and a beverage. There are 3 meats (hot dog, hamburger, and cheeseburger), 3 snacks (chips, cookies, candy), and 4 beverages (water, juice, sports drink, and soda). How many different combos exist?

jimthompson5910 (jim_thompson5910):

number of combos = (# of meats)*(# of snacks)*(# of drinks) number of combos = 3*3*4 number of combos = 9*4 number of combos = 36

OpenStudy (anonymous):

Aha! That's what I got, too! Thanks again! :) You're a whiz bomb at math!

jimthompson5910 (jim_thompson5910):

that's great, it sounds like you know what you're doing then

OpenStudy (anonymous):

Kind of - I'm taking Algebra 2, so it can sometimes be confusing. Tonight is one of those "My brain isn't working" nights. :)

jimthompson5910 (jim_thompson5910):

Definitely been there before lol

OpenStudy (anonymous):

Ha, ha - I think we all have :)